OpenStudy (anonymous):

Hey guys! just wondering, how would you divide 800, a bigger number into 500, a smaller number.

OpenStudy (lgbasallote):

bye canceling some zeroes..

OpenStudy (lgbasallote):

by*

OpenStudy (anonymous):

so, you mean like 800/500.0?

OpenStudy (radar):

by using a calculator

OpenStudy (lgbasallote):

no no no... 800/500...you can cancel 2 zeros because they both have two..so it becomes 8/5

OpenStudy (anonymous):

They make you dumb, that is why I tried solving the procedure for 4 hours.

OpenStudy (anonymous):

Lgbasalloote is 100 correct. And it checks out. Done on a calc 800/500= 1.6. 8/5=1.6

OpenStudy (radar):

500/800=5/8=0.625

OpenStudy (anonymous):

No, I mean 800 into 500.